a rectangle has sides with a length of 5cm and 12 cm. what is the length of the digaonal of the rectangle
Answer by oberobic(2304)

A rectangle has 4 90-degree angles, so you can apply the Pythagorean Theorem to find the diagonal.
c^2 = a^2 + b^2
a = 5 cm
b = 12 cm
.
c^2 = 25 + 144 = 169 cm^2
.
c = sqrt (169) = + or - 13 cm
.
A negative diagonal is spurious, so we can say:
.
Diagonal = 13 cm.
.
Done.